OpenStudy (anonymous):

A class consists of 3 boys and 6 girls willing to form 3 groups of 3 called Group A, B, C. How many ways are there to create groups such that every group includes 1 boy? How many ways are there to create groups such that all three boys are in the same group? How many ways are there to form groups such that one group has no boys?

OpenStudy (anonymous):

@Zarkon

OpenStudy (zarkon):

show me what you have so far and I'll tell you if it is correct

OpenStudy (anonymous):

For the first part, my idea is to count the number of ways to partition girls into 3 group of 2, then multiply by 3! for the number of ways the boys can fill the spaces. I'm not sure.

OpenStudy (anonymous):

\[\frac{ 6! }{ 2!2!2! } * 3!\]

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@hartnn

OpenStudy (zarkon):

I think you might be double counting

OpenStudy (anonymous):

Can you show where?

OpenStudy (zarkon):

(lets simplify the problem just a little) how many ways would you have if you were making 3 groups of 2...having 3 boys and 3 girls (again each group containing 1boy)

OpenStudy (anonymous):

Meaning each group contains 1 boy and 1 girl.

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Number of ways of choosing 3 girls into 3 groups = 3! Multiples by the number of ways of choosing 3 boys into 3 groups = 3! (3!)*(3!)

OpenStudy (zarkon):

b1g1,b2g2,b3g3 b1g1,b2g3,b3g2 b1g2,b2g1,b3g3 b1g2,b2g3,b3g2 b1g3,b2g1,b3g2 b1g3,b2g2,b3g1 did i miss any

OpenStudy (zarkon):

b1g2,b2g3,b3g2 should be b1g2,b2g3,b3g1

OpenStudy (zarkon):

you are saying then that there are 3!*3!=36...which are we missing?

OpenStudy (zarkon):

oh...if you are naming the groups ...then you are correct...if the groups are indistinguishable then you would need a different answer

OpenStudy (anonymous):

Yes the groups are named.

OpenStudy (zarkon):

then your answer to the first one is correct

OpenStudy (anonymous):

Are you very sure?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

The second part is where it all falls apart in my brain.

OpenStudy (anonymous):

Thank you though.

OpenStudy (anonymous):

Here's my idea:

OpenStudy (anonymous):

if all three boys are in the same group, find the ways of putting the girls into 2 groups of 3, then find the number of ways of putting the boys as a whole into the last group.

OpenStudy (anonymous):

@Zarkon

OpenStudy (zarkon):

the answer will look a lot like your first answer

OpenStudy (anonymous):

Can I get a hint?

OpenStudy (zarkon):

I would first determine how many ways to get the girls into the first 2 groups and the boys in the last group

OpenStudy (anonymous):

\[\frac{ 6! }{ 3!*3! } * Some factor for boys\]

OpenStudy (zarkon):

there is only one way to select the three boys for a group of 3 \[{3\choose 3}\]

OpenStudy (anonymous):

Yes, I got that fact, but the group of boys can take 3 different names. So *3?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Is it 3 or 3!?

OpenStudy (zarkon):

with 3! you would be double counting (because of the girls)

OpenStudy (anonymous):

Even if the groups have names?

OpenStudy (zarkon):

when in doubt reduce the problem so something simple what if you had 2 girls and 1 boy and you made 3 groups of 1 then A B C g1,g2,b g2,g1,b g1,b,g2 g2,b,g1 b,g1,g2 b,g2,g1 the number of ways is \[\frac{2!}{1!1!}\times 3=6\]

OpenStudy (anonymous):

Okay, thanks. Now the last part.

OpenStudy (anonymous):

To have exactly 1 with no boys, some group must have 2 boys while another must not have 3 boys.

OpenStudy (zarkon):

if it is exactly one boy then one group will have none, one group will have 1 and one group will have 2

OpenStudy (anonymous):

Yes, but I want to do it using the result from the previos part so it'll be quicker.

OpenStudy (anonymous):

So, my idea is: n(1 group with none) - n(One group with 3)

OpenStudy (zarkon):

total-n(1 group with 3 boys)-N(each group with exactly 1)

OpenStudy (anonymous):

Yeah, now I see where I flunked. I knew there was a relationship. Are you very sure though?

OpenStudy (zarkon):

I did it both ways...direct and with the 'complement' and I get the same answer

OpenStudy (anonymous):

And what is total? Can you share your answers?

OpenStudy (zarkon):

Look at \[\frac{6!}{1!3!2!}\cdot\frac{3!}{2!1!}\cdot 3!\] and \[\frac{9!}{3!3!3!}-\frac{6!}{3!3!}\cdot 3-\frac{6!}{2!2!2!}\cdot 3!\]

OpenStudy (zarkon):

can you parse what I wrote?

OpenStudy (zarkon):

do you understand?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Wait, define the first part.

OpenStudy (zarkon):

the \(\dfrac{6!}{1!3!2!}\) is the number of ways to arrange the girls the \(\dfrac{3!}{2!1!}\) is the number of ways to arrange the boys and the \(3!\) is the number of ways to arrange these into the 3 labeled groups.

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

Yes, I'm just verifying the answer. Give me a minute please.

OpenStudy (anonymous):

I'm not getting the same answer.

OpenStudy (zarkon):

you should

OpenStudy (anonymous):

Total is 9!/(3!)^3 right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Check to make sure both are the same. One aboyt twice the other.

OpenStudy (zarkon):

they are both 1080

OpenStudy (zarkon):

jas

OpenStudy (zarkon):

they are both 1080... type 9!/(3!^3)-(6!/((3!)^2)*3+6!/(2!)^3*3!) and 6!/(1!*3!*2!)*3!/(2!*1!)*3! into wolfram alpha

OpenStudy (zarkon):

or just copy and paste

OpenStudy (zarkon):

I have to go...I'll be back later to see how things went.

OpenStudy (anonymous):

Okay, thanks a lot.

OpenStudy (anonymous):

@Zarkon I have a question about your solution.